OpenStudy (anonymous):

Among all the unit vectors u= [x y z] find the one for which the sum of x+2y+5z is minimal

OpenStudy (turingtest):

um... how about \(\vec e=\langle 1,0,0\rangle\) (not sure I understand the question)

OpenStudy (anonymous):

This will do it \[\left(-\frac{1}{\sqrt{30}},-\sqrt{\frac{2}{15}},-\sqrt{\frac{5}{6}}\right) \] The minimum of x+2y+5z will be about -5.47723. One is tempted to take (0,0,-1), but it is not right. It is easy to work with spherical coordinates. All unit vectors look like \[(\cos (\theta ) \sin (\phi ),\sin (\theta ) \sin (\phi ),\cos (\phi )),\quad 0\le \theta \le 2\pi, 0\le \phi \le \pi \] try to minimize the quantity\[ x+2y+5z =2 \sin (\theta ) \sin (\phi )+\cos (\theta ) \sin (\phi )+5 \cos (\phi ) \] I leave the details for you.

OpenStudy (zarkon):

I would just use Lagrange multipliers on the original equations/function

OpenStudy (anonymous):

You can also do that. Probably do it and show the asker the details.

OpenStudy (anonymous):

thank you ill work on getting the details